A license plate has 6 positions for letters and numbers. The first two positions must be letters. The last four positions must be nonzero whole numbers. How many different license plates can be produced using this combination of letters and numbers?
Found 2 solutions by stanbon, Fombitz:
Answer by stanbon(48569)

A license plate has 6 positions for letters and numbers. The first two positions must be letters. The last four positions must be nonzero whole numbers. How many different license plates can be produced using this combination of letters and numbers?
------------------
# = 26^2*9^4 = 4435236
==============
Cheers,
Stan H.

Answer by Fombitz(13823)

License plates that run from AA1111 to ZZ9999
Position one - 26 (A-Z)
Position two - 26 (A-Z)
Position three - 9 (1-9)
Position four - 9 (1-9)
Position five - 9 (1-9)
Position six - 9 (1-9)
Total number of plates available = 26*26*9*9*9*9 = 4,435,236